OpenStudy (anonymous):

Solve the quadratic equation x^2 - 56 = 10x by factoring. A. {4, 14} B. {-14, 4} C. {-14, -4} D. {-4, 14}

OpenStudy (solomonzelman):

\[ x^2 - 56 = 10x\]\[ x^2 - 56 \color{red} {-10x}= 10x\color{red} {-10x}\]\[x^2\color{red} {-10x}-56=0\] \[(x-14)(x+4)=0\]\[\color{red} {you~~can~~do~~i t~~from~~here.}\]

OpenStudy (anonymous):

Thanks

OpenStudy (solomonzelman):

Anytime!